OpenStudy (anonymous):

The table shows movie preferences among high school and college students: Movie Preference Action Drama Comedy Row totals High School 40 23 12 75 College 20 55 28 103 Column totals 60 78 40 178 Which of the following are possible positive associations or trends? For the condition of type of movie, high school students prefer action—40%. For the condition of type of movie, college students prefer drama—70.5%. For the condition of type of movie, high school students prefer action—67%. For the condition of type of movie, college students prefer drama—55%.

OpenStudy (mayankdevnani):

@blackwonder just find out percentage and get your answer

OpenStudy (anonymous):

how do i find the percentage

OpenStudy (mayankdevnani):

Remember :- \[\huge \bf percentage=\frac{no~.of~ items~ used}{total~no.~of~items}\]

OpenStudy (mayankdevnani):

then multiply by 100,you get percentage. percentage is denoted by "%"

OpenStudy (mayankdevnani):

i can solve your question and if you have any doubt in my any step,please comment on it

OpenStudy (anonymous):

so am i just fining the average and multiplying it by 100

OpenStudy (mayankdevnani):

Condition type of movie for high school :- \[\large \bf percentage=\frac{no.~of~items~used}{total~no.~of~items}=\frac{40+23+12}{60+78+40}\] \[\large \bf =\frac{75}{178}\] then multiply by 100,we get \[\large \bf \frac{75}{178} \times 100=42 percent =40 percent\]

OpenStudy (mayankdevnani):

so your option A is correct

OpenStudy (mayankdevnani):

now similarly solve your question for college students.

OpenStudy (mayankdevnani):

getting it ? @blackwonder

OpenStudy (anonymous):

ya thanks so much man this is my first time using this site so i dont understand the whole medal thing how does that work

OpenStudy (mayankdevnani):

welcome :) Just click BEST RESPONSE just right side to my comment. This pic illustrates you :-

OpenStudy (mayankdevnani):

okk. @blackwonder

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks for thhe help

OpenStudy (mayankdevnani):

welcome :)